OpenStudy (anonymous):

Help solving cos^2 (2x) -sin^2 (2x)=0 and sin(2x)=sin x ?

OpenStudy (arindameducationusc):

Do you have options? i guess I found the first answer... but need to check if it matches @Kiritina

OpenStudy (anonymous):

cos^2 (2x) -sin^2 (2x) = cos(4x)

OpenStudy (anonymous):

No answers available, sorry. I'm not sure where to start with either equation.

OpenStudy (arindameducationusc):

can you manage now @Kiritina or should i do the solution?

OpenStudy (anonymous):

I'd love it if you helped with the solution!

OpenStudy (jhannybean):

lets see... \[\cos^2(2x) -\sin^2(2x) = 0\]\[\color{red}{\sin(2x) = \sin(x)}\] is the red portion what you're given? Do you mind posting the actual problem by a picture?

OpenStudy (anonymous):

https://gyazo.com/976ec9f78a10f5487c8421aaa06e36a0

OpenStudy (arindameducationusc):

O yes @Jhannybean is right I just solved the black portion I didn't think about the red!

OpenStudy (arindameducationusc):

ok both are different problems hehe :D

OpenStudy (jhannybean):

So you're looking for a general solution... x_x I'm terrible at finding those.

OpenStudy (arindameducationusc):

okay 1st question we have cos^2x-sin^2x=0 =>cos4x=0 \[=>4x=\cos^{-1} (0)\] \[=\cos^{-1} (\cos (\frac{ \pi }{ 2}))\] =>4x=pi/2 >x=pi/8

OpenStudy (arindameducationusc):

2nd question, sin2x=2sinxcosx so, sin2x=sinx =>2sinxcox/sinx=1 =>2cosx=1 =>cosx=1/2 x=cos inverse(1/2) x=cos inverse (cos pi/3) x=pi/3

OpenStudy (arindameducationusc):

@Kiritina if you liked my response, dont forget to medal and fan. Thank you